# ObjectRelator: Enabling Cross-View Object Relation Understanding in Ego-Centric and Exo-Centric Videos

### v1

@@ Abstract
In this paper, we focus on the Ego-Exo Object Correspondence task, an emerging challenge in the field of computer vision that aims to map objects across ego-centric and exo-centric views. We introduce ObjectRelator, a novel method designed to tackle this task, featuring two new modules: Multimodal Condition Fusion (MCFuse) and SSL-based Cross-View Object Alignment (XObjAlign). MCFuse effectively fuses language and visual conditions to enhance target object localization, while XObjAlign enforces consistency in object representations across views through a self-supervised alignment strategy. Extensive experiments demonstrate the effectiveness of ObjectRelator, achieving state-of-the-art performance on Ego2Exo and Exo2Ego tasks with minimal additional parameters. This work provides a foundation for future research in comprehensive cross-view object relation understanding highlighting the potential of leveraging multimodal guidance and cross-view alignment. Codes and models will be released to advance further research in this direction.
@@ 1 Introduction
Advancements in video applications, including classification [21, 62, 41], captioning [25, 15, 64], and generation [57, 54, 61], have largely focused on exo-centric (third-person) videos. In contrast, progress on ego-centric (first-person) videos, crucial for fields like virtual reality and robotics, are relatively behind. Essentially, most efforts have concentrated on developing ego-centric models [24, 4, 33, 9, 65, 58] to bridge ego- and exo-centric video understanding, while fewer studies [48, 55, 39] have explored the relationships between the two views directly, due to the lack of cross-view datasets. The recent release of the Ego-Exo4D dataset [10], a large-scale, temporally aligned ego-exo dataset, enables the study of cross-view object relation understanding, offering a promising approach to bridging the ego-exo gap.
@@
Specifically, in this paper, we tackle the task of Ego-Exo Object Correspondence [^1] Originally termed “Ego-Exo Correspondence” in Ego-Exo4D [10]. Renamed for clarity, emphasizing object-level mapping., first introduced in Ego-Exo4D [10]. As shown in Fig.1, given object queries from one perspective (e.g., ego view), the task involves predicting the corresponding object masks in another perspective (e.g., exo view). Unlike prior segmentation tasks [12, 22, 27, 46, 69], which use categories, text descriptions, or visual hints as prompts, this task focuses on cross-view object understanding, utilizing visual cues from one perspective to predict the corresponding objects in a completely different perspective. This makes the direct application of existing methods nontrivial. Critically, few studies have explored ego-exo object correspondence, with PSALM [68] standing out as a notable exception. Particularly, PSALM [68] integrates Mask2Former [1] with large language models (LLMs), outperforming flagship segmentation models e.g., LISA [27], SEEM [69], and SAM [23], while also enabling zero-shot learning (ZSL) on the ego-exo object correspondence task.
However, despite its state-of-the-art performance on other segmentation tasks, PSALM faces challenges with the ego-exo object correspondence task due to two key limitations: 1) Relying solely on object masks underutilizes semantic information, limiting the LLM’s language understanding. For instance, it may segment objects with similar shapes to the query but incorrect categories. 2) PSALM does not account for the cross-view relationships between objects in ego-exo scenarios, causing it to fail when significant visual appearance shifts are observed. These challenges raise two critical questions: ① Can incorporating language conditions with mask queries improve object localization? ② Can we ensure consistency across perspectives for the same objects?
@@
To address these key challenges, we propose two innovative modules. First, the Multimodal Condition Fusion (MCFuse) module is proposed to tackle the question ①. While the task provides only object query masks, MCFuse enhances this by incorporating semantic information from large vision-language models like LLaVA [35], which generates language descriptions of the query object. We then combine both the object mask and the generated text in the embedding space using a lightweight cross-attention mechanism with residual connection and learnable fusion weight. This fusion enables the model to leverage both visual and linguistic cues, leading to more accurate localization of the target object. Next, to deal with question ②, we introduce the Cross-View Object Alignment (XObjAlign) module. This module aligns paired ego and exo object masks into a shared latent space, enforcing proximity between them. Utilizing an effective self-supervised learning approach, we optimize the distance similarity between the masks, which helps the model better recognize objects across different perspectives and overcome challenges posed by view changes.
[…]
We summarize our contributions as, 1) Ego-Exo Object Correspondence Task: Early exploration of the challenging task is performed highlighting its unique difficulties from traditional segmentation. 2) Multimodal Condition Fusion (MCFuse): MCFuse enhances object localization by effectively combining visual and linguistic cues, using language descriptions from vision-language models. 3) Cross-View Object Alignment (XObjAlign): XObjAlign ensures consistency across ego and exo perspectives, improving the model’s robustness to view changes. 4) ObjectRelator Framework: By integrating MCFuse and XObjAlign, we develop the novel ObjectRelator, achieving state-of-the-art results on both Ego2Exo and Exo2Ego tasks. 5) Flexibility and Adaptability: We demonstrate ObjectRelator’s flexibility via joint training and single-condition testing.
@@ 2 Related Work
Ego-Exo Video Understanding. Both ego- and exo-centric video perspectives are crucial for understanding the world. Extensive work has been done in single-view video understanding: For ego-centric videos, numerous methods address tasks such as video classification [21, 62, 8, 41], captioning [25, 15, 64], video generation [57, 54, 61]. While ego-centric research has historically lagged behind exo-centric, recent efforts including datasets such as Ego4D [9], EK100 [4], HoloAssist [58], AssistQ [60], as well as a variety of methods [32, 24, 7, 19, 33, 65, 43, 66] have brought substantial progress. However, far fewer studies directly address the relationship between ego and exo views. Notable exceptions have ego-exo translation/generation [55, 39, 2, 38, 36]. One factor leading to this gap is that existing datasets are often small in scale or lack comprehensive annotations [5, 53, 18, 44, 26, 14, 50]. Recently, the release of Ego-Exo4D [10], a large-scale, richly annotated, and time-synchronized multimodal dataset, has opened new chances for advancing ego-exo research. In this paper, we focus on the ego-exo object correspondence task proposed in Ego-Exo4D, aiming to establish a strong starting point with our approach.
@@
UNINEXT [63], SEEM [69], PSALM [68] Universal Segmentation Text/Visual Prompts, Prompt Decoder with referring to image
ObjectRelator (Ours) Ego-Exo Object Correspondence Text+Visual Prompts, Prompt Decoder with referring to image -
@@
Segmentation Models. Segmentation models can roughly be grouped into several main categories based on their input types: generic segmentation, referring segmentation, and interactive segmentation. We refer to generic segmentation [40] as those more classical segmentation tasks e.g., semantic segmentation [28], instance segmentation [11], and panoptic segmentation [22]. Generic segmentation methods don’t need extra text prompts or visual prompts, for example, Mask-RCNN [12] and Mask2Former [1]. In contrast, referring segmentation [17] requires textual descriptions to guide the segmentation. Flagship examples include LISA [27], GLAMM [46], PixelLM [49]. The interactive segmentation [45] utilizes visual hints or prompts such as bounding boxes, key points, or masks provided within the image to guide the segmentation results. Well-known examples have OMG-Seg [30], SAM [23], SAM2 [47]. The latest advancements such as UNINEXT [63], SEEM [69], PSALM [68] accept multiple conditions as input, which we termed as “Universal Segmentation”. However, none of them are specifically tailored for the ego-exo object correspondence, which is essentially challenged by the huge view change between the first-person and third-person observations. Thus, how to build a good segmentation model for this new task still remains uncovered.
Differences of methods are summarized in Tab.1. Typically, SAM [23], OMG-Seg [30], and SAM2 [47] decode visual prompts solely from hints such as points, masks, and bounding boxes, without referencing the corresponding image. These approaches require strict alignment between the mask and testing image, limiting adaptability for tasks like ours, where object locations and shapes vary inherently. By contrast, UNINEXT [63], SEEM [69], and PSALM [68] decode visual prompts with image reference, making them more comparable to our method. Importantly, we differ from existing segmentation approaches by focusing on cross-view challenges and exploring the fusion of multiple conditions from a single image.
@@ 3 Methodology
Task Formulation. The Ego-Exo Object Correspondence involves mapping objects between ego- and exo-centric video perspectives as in Fig. 1. Formally, given a query video $V_{\text{query}}$ (e.g., an ego-centric or exo-centric video) and a target video $V_{\text{target}}$ (e.g., an exo-centric or ego-centric video), along with object masks $\mathbf{M}_{\text{query}}$ indicating the object of interest in $V_{\text{query}}$, the objective is to predict the corresponding object masks $\mathbf{M}_{\text{target}}$ in $V_{\text{target}}$. The query and target videos are assumed to be temporally aligned, ensuring that corresponding frames represent the same time points. Additionally, each video pair may contain multiple objects of interest. Any semantic information e.g., the categories of the objects are not provided in the benchmark. Models are expected to leverage cross-view relationships to locate and segment the corresponding objects in the target perspective.
To structure and simplify usage, we rename the two sub-tasks as Ego2Exo, where the ego-centric video serves as the query and the exo-centric video as the target, and Exo2Ego, where the exo-centric video is the query and the ego-centric video is the target.
PSALM Baseline. As discussed in Sec. 2 and shown in Tab. 1, the UNINEXT [63], SEEM [69], PSALM [68] could be considered as the baselines. Among them, we select PSALM for its: 1) validated superior performance over UNINEXT and SEEM models on several various benchmarks; 2) demonstrated zero-shot learning (ZSL) transfer ability to our task. Particularly, PSALM combines Mask2Former [1], the potent generic segmentation model, with a powerful LLM leveraging the strengths of both. Briefly, PSALM first takes an image, an instruction prompt, a condition prompt (e.g., categories, text, or visual hint), and mask tokens as input to the LLM resulting in the condition embedding, mask embedding. It then predicts masks in a Mask2Former style based on the image and these LLM extracted embeddings. The model is trained by minimizing the loss, $\mathcal{L}_{mask}$, which is calculated by comparing the predicted masks to the ground truth.
Figure 2: Overview of ObjectRelator. Ego2Exo is used as an example. The dashed lines mean the base PSALM relies solely on the ego mask prompt as the condition. The colorful lines represent our multimodal condition fusion and cross-view object alignment.
@@ 3.1 ObjectRelator Framework
Our proposed ObjectRelator operates at the frame level, focusing on studying object relations between ego-exo frames, leaving the exploration of temporal information as future work. The framework is illustrated in Fig. 2. Overall, we retain the core PSALM modules: Mask Token Extraction, Visual Encoder, MM (Multimodal) Projector, LLM, Pixel Decoder, Mask Generator, and also its loss function $\mathcal{L}_{mask}$, while introducing two novel components, Multimodal Condition Fusion (MCFuse) and Cross-View Object Alignment (XObjAlign), highlighted in orange and light green colors.
@@
For clarity, we use Ego2Exo as an example: each training step involves a synchronized ego image as the query and an exo image as the target. For each object in the paired data, we denote the ego image, ego object mask, exo image, and exo object mask as $I^{ego}$, $m^{ego}$, $I^{exo}$, and $m^{exo}$ (ground truth), respectively. The following steps are:
@@
1) Preparing Inputs for LLM: For the target exo image $I^{exo}$, we use the Visual Encoder, MLP Projector, and Pixel Decoder to extract features $f_{I}^{exo}$ and $f_{I}^{{exo}_{multi}}$, where $f_{I}^{{exo}_{multi}}$ indicates the feature has multiple layers. Unlike the ZSL PSALM model, which only uses the single ego object mask as a prompt for exo, we leverage both the visual object region and a language description as prompts. Thus, our instruction prompt $P_{ins}$ is formulated as: “This is an image. Please segment by given regions and instruction.”, where the region hint is ego mask prompt $P_{V}^{ego}$ and the instruction hint is the text instruction prompt $P_{T}^{ego}$. Specifically, $P_{V}^{ego}$ is formed by combing the ego image $I^{ego}$ and its object mask $m^{ego}$, while $P_{T}^{ego}$ describes the masked object in $I^{ego}$, generated via the pre-trained multimodal model LLaVA [35]. Additionally, to enforce consistency between ego and exo objects, we use the ground truth $m^{exo}$ to form the exo mask prompt $P_{V}^{exo}$. Note that although named a prompt, $P_{V}^{exo}$ is not used for mask generation but only for the XObjAlign module. We then extract the instruction tokens $T_{ins}$, ego text condition token $T_{T}^{ego}$, ego visual condition token $T_{V}^{ego}$, exo visual token $T_{V}^{exo}$, and mask token $T_{M}$, following the same token extraction method e.g., the Mask Token Extraction module as the base PSALM.
@@
2) Forwarding the Network: We process image features $f^{exo}_{I}$, $f^{{exo}_{multi}}_{I}$, instruction tokens $T_{ins}$, ego text tokens $T_{T}^{ego}$, ego visual tokens $T_{V}^{ego}$, exo visual tokens $T_{V}^{exo}$, and mask tokens $T_{M}$. To feed these into the pretrained LLM, we create two inputs: one concatenating $f^{exo}_{I}$, $T_{ins}$, $T_{T}^{ego}$, $T_{V}^{ego}$, and $T_{M}$, and the other with $f^{exo}_{I}$, $T_{ins}$, $T_{T}^{ego}$, $T_{V}^{exo}$, and $T_{M}$. Feeding both inputs into the LLM yields ego text emb $E_{T}^{ego}$, ego visual emb $E_{V}^{ego}$, exo visual emb $E_{V}^{exo}$, and mask emb $E_{M}$. The embeddings $E_{V}^{ego}$ and $E_{V}^{exo}$, representing the same object from different views, are passed through the XObjAlign module to compute the cross-object consistency loss $\mathcal{L}_{X{obj}}$. Meanwhile, $E_{T}^{ego}$ and $E_{V}^{ego}$ are sent to the MCFuse module to generate the fused multimodal ego embedding $E_{MM}^{ego}$. Finally, the multiscale image feature $f^{{exo}_{multi}}_{I}$, the fused ego multimodal emb $E_{MM}^{ego}$, and the mask emb $E_{M}$ are passed to the Mask Generator to predict the exo masks and compute the loss $\mathcal{L}_{mask}$.
@@
3) Training and Inference: Our model has two training stages: a) S1: The MCFuse module is trained with the $\mathcal{L}_{mask}$ loss to initialize it well for the second stage. b) S2: All modules (indicated by the fire icon) are trained jointly with the losses $\mathcal{L}_{mask}$ and $\mathcal{L}_{X{obj}}$. During inference, the XObjAlign module and also exo mask prompt $P^{exo}_{V}$ are removed, and the remaining model is used for inference.
3.2 Key Module: Multimodal Condition Fusion

### v2

# ObjectRelator: Enabling Cross-View Object Relation Understanding Across Ego-Centric and Exo-Centric Perspectives

@@ Abstract
Bridging the gap between ego-centric and exo-centric views has been a long-standing question in computer vision. In this paper, we focus on the emerging Ego-Exo object correspondence task, which aims to understand object relations across ego-exo perspectives through segmentation. While numerous segmentation models have been proposed, most operate on a single image (view), making them impractical for cross-view scenarios. PSALM [75], a recently proposed segmentation method, stands out as a notable exception with its demonstrated zero-shot ability on this task. However, due to the drastic viewpoint change between ego and exo, PSALM fails to accurately locate and segment objects, especially in complex backgrounds or when object appearances change significantly. To address these issues, we propose ObjectRelator, a novel approach featuring two key modules: Multimodal Condition Fusion (MCFuse) and SSL-based Cross-View Object Alignment (XObjAlign). MCFuse introduces language as an additional cue, integrating both visual masks and textual descriptions to improve object localization and prevent incorrect associations. XObjAlign enforces cross-view consistency through self-supervised alignment, enhancing robustness to object appearance variations. Extensive experiments demonstrate ObjectRelator’s effectiveness on the large-scale Ego-Exo4D benchmark and HANDAL-X (an adapted dataset for cross-view segmentation) with state-of-the-art performance. Code is available at: http://yuqianfu.com/ObjectRelator.
@@ 1 Introduction
Advancements in vision tasks [26, 69, 46, 30, 20, 71] have largely focused on exo-centric (third-person) perspectives. In contrast, progress on ego-centric (first-person) perception is relatively behind. To bridge this gap, most efforts have focused on developing ego-centric models [29, 6, 38, 12, 72, 65] while fewer studies [53, 62, 54, 44] have directly explored the relationships between the two views, primarily due to the lack of cross-view datasets. The recent release of the Ego-Exo4D [13], a large-scale, temporally aligned ego-exo dataset, enables the study of cross-view object relation understanding, offering a promising new direction.
@@
Specifically, in this paper, we tackle the task of Ego-Exo Object Correspondence [^1] Originally termed “Ego-Exo Correspondence” in Ego-Exo4D [13]. Renamed for clarity, emphasizing object-level mapping., first introduced in Ego-Exo4D [13]. As shown in Fig.1, given object queries from one perspective (e.g., ego view), the task involves predicting the corresponding object masks in another perspective (e.g., exo view). Solving this task unlocks new possibilities in VR and robotics, e.g., enabling virtual agents or robots to manipulate ego-view actions by learning from exo-view demonstrations. However, despite the existence of numerous segmentation models, this task remains non-trivial: most prior approaches operate on a single image (view), whereas Ego-Exo Object Correspondence requires leveraging visual cues from one perspective to identify corresponding objects in a completely different one.
Critically, PSALM [75], a recently proposed vision-language-based segmentation method, stands out as a notable exception. Combining Mask2Former [2] with the powerful Large language model (LLM), PSALM outperforms flagship segmentation models such as LISA [32] and SEEM [77], while also enabling zero-shot learning (ZSL) for the ego-exo object correspondence task. However, despite achieving state-of-the-art performance on other segmentation tasks, PSALM struggles with the unique challenges of ego-exo object correspondence: i) Complex backgrounds in the exo view often introduce distractors, causing the model to segment objects with similar shapes but incorrect categories; ii) Significant variations in object appearance between ego and exo views further increase the task’s difficulty. These challenges motivate us to enhance PSALM by: ① improving object localization and ② ensuring object consistency across perspectives.
@@
To this end, we propose two innovative modules, i.e., Multimodal Condition Fusion (MCFuse) and Cross-View Object Alignment (XObjAlign). MCFuse addresses ① by incorporating language descriptions into the task. While the benchmark provides only object query masks, MCFuse managed to generate textual descriptions of the query object via large vision-language models like LLaVA [40]. Then, a lightweight cross-attention with residual connections and learnable fusion weights fuses the visual mask and generated text in the embedding space. This enables the model to leverage both visual and linguistic cues, improving target object localization. To tackle ②, XObjAlign is proposed to align paired ego and exo objects in a shared latent space, enforcing proximity between them. Employing a self-supervised learning approach, we optimize the similarity between ego and exo representations, particularly at the object level, improving the model’s ability to recognize objects across different perspectives and address challenges arising from viewpoint variations.
By integrating the proposed MCFuse and XObjAlign into PSALM, we propose our ObjectRelator. We validate its effectiveness through extensive experiments on the large-scale Ego-Exo4D benchmark, achieving state-of-the-art (SOTA) performance with significant improvements over existing competitors. Additionally, we introduce HANDAL-X, an adapted cross-view object segmentation dataset with robotics-ready, manipulable objects as a new testbed, and demonstrate our method’s superiority.
We summarize our contributions as:
$\bullet$ Toward Ego-Exo Object Correspondence Task: We conduct an early exploration of this challenging task, analyzing its unique difficulties, constructing several baselines, and proposing a new method.
$\bullet$ ObjectRelator Framework: We introduce ObjectRelator, a cross-view object segmentation method combining MCFuse and XObjAlign. MCFuse for the first time introduces the text modality into this task and improves localization using multimodal cues for the same object(s), while XObjAlign boosts performance under appearance variations with an object-level consistency constraint.
$\bullet$ New Testbed &\&& SOTA Results: Alongside Ego-Exo4D, we present HANDAL-X as an additional benchmark. Our proposed ObjectRelator achieves state-of-the-art (SOTA) results on both datasets.
@@ 2 Related Work
Ego-Exo Perception Understanding. Both ego- and exo-centric perceptions are crucial for understanding the world. Extensive work has been done in single-view understanding: For exo-centric view, numerous methods address tasks such as video classification [26, 69, 10, 11, 46], captioning [30, 20, 71], detection and segmentation [61, 5, 1, 76], and generation [64, 60, 68, 18]. While ego-centric research has historically lagged behind exo-centric, recent efforts including datasets such as Ego4D [12], EK100 [6], HoloAssist [65], AssistQ [67], as well as a variety of methods [37, 29, 9, 24, 38, 72, 48, 73] have brought substantial progress. However, far fewer studies directly address the ego-exo relationships. Notable exceptions have ego-exo translation/generation [62, 44, 3, 43, 41]. One reason for this gap is that existing datasets are often small or lack comprehensive annotations [7, 59, 23, 49, 31, 19, 56]. The recent release of Ego-Exo4D [13], a large-scale, richly annotated, and time-synchronized dataset, opens new chances for advancing ego-exo research. In this paper, we focus on the ego-exo object correspondence task proposed in Ego-Exo4D, aiming to establish a good starting point.
@@
UNINEXT [70], SEEM [77], PSALM [75] Universal Segmentation Text or Visual Prompts, Prompt Decoder with referring to image
ObjectRelator (Ours) Ego-Exo Object Correspondence Visual Prompts, Prompt Decoder with referring to image -
@@
Segmentation Models. Segmentation models can roughly be grouped into several categories based on input types: we refer to generic segmentation [45] as those more classical segmentation tasks e.g., semantic segmentation [33], instance segmentation [15], and panoptic segmentation [27]. Generic segmentation methods (e.g., Mask R-CNN [16] and Mask2Former [2]) don’t need extra text prompts or visual prompts. In contrast, referring segmentation [22] requires textual descriptions to guide the segmentation. Flagship examples include LISA [32], GLAMM [51], PixelLM [55]. Interactive segmentation [50] utilizes visual hints such as bounding boxes, key points, or masks provided within the image as prompts. Well-known examples have SAM [28], OMG-Seg [35], SAM2 [52]. The latest advancements such as UNINEXT [70], SEEM [77], PSALM [75] accept multiple conditions as input, which we termed as universal segmentation. However, none of them are specifically tailored for the ego-exo object correspondence, which is essentially challenged by the huge view change between two images.
More differences of methods are summarized in Tab.1. Typically, SAM [28], OMG-Seg [35], and SAM2 [52] decode visual prompts solely from hints such as points, masks, and bounding boxes, without referencing the corresponding image. These approaches require strict alignment between the mask and testing image, limiting adaptability for tasks like ours, where object locations and shapes vary inherently. By contrast, UNINEXT [70], SEEM [77], and PSALM [75] decode visual prompts with image reference, making them more comparable to us. Importantly, we differ from existing segmentation approaches by focusing on cross-view challenges and exploring the fusion of multiple conditions.
@@ 3 Methodology
Task Formulation. The Ego-Exo Object Correspondence involves mapping objects between ego-exo perspectives as in Fig. 1. Formally, given a query view $I^{*}/V^{*}$ (e.g., an ego-centric or exo-centric image/video) and a target view $I/V$ (e.g., an exo-centric or ego-centric image/video), along with object mask(s) $m^{*}$ indicating the object of interest in query view, the objective is to predict the corresponding object mask(s) in the target view $m$. The query and target images/videos are assumed to be temporally aligned, ensuring that corresponding frames represent the same time points. Additionally, each pair may contain multiple objects of interest. Any semantic information (e.g., the categories of the objects) is not originally provided in the benchmark.
To structure and simplify usage, we rename the two sub-tasks as Ego2Exo, where the ego-centric view serves as the query and the exo-centric view as the target, and Exo2Ego, where the roles of ego and exo are reversed.
PSALM Baseline. As discussed in Sec. 2 and shown in Tab. 1, the UNINEXT [70], SEEM [77], PSALM [75] could be considered as the baselines. Among them, we select PSALM for its: 1) validated superior performance over UNINEXT and SEEM models on several various benchmarks; 2) demonstrated zero-shot learning (ZSL) transfer ability to our task. Particularly, PSALM combines Mask2Former [2], the potent generic segmentation model, with a powerful LLM leveraging the strengths of both.
Briefly, PSALM mainly contains a Visual Encoder, MM (Multimodal) Projector, LLM, Pixel Decoder, Mask Generator. For each segmentation, PSALM takes an image, an instruction prompt, a condition prompt (e.g., categories, text, or visual hints), and mask tokens as input, feeds their corresponding feature or tokens into the LLM resulting in the condition embedding, mask embedding. After that,  PSALM predicts masks in a Mask2Former style: it inputs the multiscale image feature, condition embedding, and mask embedding to the Mask Generator. The image feature for LLM is extracted by the Visual Encoder and MM Projector, while the multiscale image feature for the Mask Generator is extracted by the Visual Encoder and Pixel Decoder. During the training, loss $\mathcal{L}_{mask}$ from Mask Generator is used to optimize all the modules except the visual encoder; During inference, the Mask Generator generates the predicted mask. For more details, e.g., token extraction method and loss function, please refer to PSALM [75].
Figure 2: Overview of ObjectRelator. Ego2Exo is used as an example. Our method builds on the PSALM baseline (pink blocks) and tailors it for Ego-Exo Object Correspondence with two novel modules: Multimodal Condition Fusion and Cross-View Object Alignment.
@@ 3.1 ObjectRelator Framework
Our ObjectRelator operates at the frame level, focusing on studying object relations between ego-exo frames, leaving the exploration of temporal information as future work. The framework is illustrated in Fig. 2. Overall, we retain the core PSALM modules: Visual Encoder, MM (Multimodal) Projector, LLM $\mathcal{H}()$, Pixel Decoder, Mask Generator $\mathcal{G}()$, loss function $\mathcal{L}_{mask}$, and also its way of extract tokens, while introducing two novel components, Multimodal Condition Fusion (MCFuse) and Cross-View Object Alignment (XObjAlign), highlighted in orange and light green colors.
@@
For clarity, we use Ego2Exo as an example: each training step involves a synchronized ego image as the query and an exo image as the target. For each object in the paired data, we denote the ego image, ego object mask, exo image, and exo object mask as $I^{*}$, $m^{*}$, $I$, and $m$ (ground truth), respectively. The following steps are:
@@
1) Preparing Inputs for LLM: For the target exo image $I$, we use the Visual Encoder, MM Projector, and Pixel Decoder to extract features $f_{I}$ and $f_{I}^{{}^{\prime\prime\prime}}$, where $f_{I}^{{}^{\prime\prime\prime}}$ indicates the feature has multiple layers. Unlike the ZSL PSALM model, which only uses the single ego object mask as a prompt for exo, we leverage both the visual object region and a language description as prompts. Thus, our instruction prompt $P_{ins}$ is formulated as: “This is an image. Please segment by given regions and instruction.”, where the region hint is ego mask prompt $P_{vis}^{*}$ and the instruction hint is the text instruction prompt $P_{txt}^{*}$. Specifically, $P_{vis}^{*}$ is formed by combing the ego image $I^{*}$ and its object mask $m^{*}$, while $P_{txt}^{*}$ describes the masked object in $I^{*}$, generated via the pre-trained multimodal model LLaVA [40]. Additionally, to enforce consistency between ego and exo objects, we use the ground truth $m$ to form the exo mask prompt $P_{vis}$. Note that although named a prompt, $P_{vis}$ is not directly used for mask generation but mainly for the XObjAlign module during training. We then extract the instruction tokens $T_{ins}$, ego text condition token $T_{txt}^{*}$, ego visual condition token $T_{vis}^{*}$, exo visual token $T_{vis}$, and mask token $T_{M}$, following the same token extraction method as the base PSALM.
2) Forwarding the Network: We process image features $f_{I}$, $f_{I}^{{}^{\prime\prime\prime}}$, instruction tokens $T_{ins}$, ego text tokens $T_{txt}^{*}$, ego visual tokens $T_{vis}^{*}$, exo visual tokens $T_{vis}$, and mask tokens $T_{M}$. To feed them into the pretrained LLM $\mathcal{H}()$, we create two inputs: one concatenating $f_{I}$, $T_{ins}$, $T_{txt}^{*}$, $T_{vis}^{*}$, and $T_{M}$, and the other with $f_{I}$, $T_{ins}$, $T_{txt}^{*}$, $T_{vis}$, and $T_{M}$. Input both to the LLM yields ego text emb $E_{txt}^{*}$, ego visual emb $E_{vis}^{*}$, exo visual emb $E_{vis}$, and mask emb $E_{M}$ as,
$E_{txt}^{*},E_{vis}^{*},E_{M}=\mathcal{H}(f_{I},T_{ins},T_{txt}^{*},T_{vis}^{*},T_{M}).$ (1)
[…]
@@
The embeddings $E_{vis}^{*}$ and $E_{vis}$, representing the same object from different views, are passed through the XObjAlign module to compute the cross-object consistency loss $\mathcal{L}_{X{obj}}$. Meanwhile, $E_{txt}^{*}$ and $E_{vis}^{*}$ are sent to the MCFuse module to generate the fused ego multimodal condition embedding $E_{con}^{*}$. Finally, the multiscale image feature $f_{I}^{{}^{\prime\prime\prime}}$, the fused ego multimodal emb $E_{con}^{*}$, and the mask emb $E_{M}$ are passed to the Mask Generator $\mathcal{G}()$ to predict the exo masks and compute the loss $\mathcal{L}_{mask}$ as,
$\mathcal{L}_{mask}=\mathcal{G}(f_{I}^{{}^{\prime\prime\prime}},E_{con}^{*},E_{M}).$ (3)
@@
3) Training and Inference: Our model has two training stages: a) S1: The MCFuse module is trained with the $\mathcal{L}_{mask}$ loss to initialize it well for the second stage. b) S2: All modules (indicated by the fire icon) are trained jointly with the final loss $\mathcal{L}=\mathcal{L}_{mask}+\mathcal{L}_{X{obj}}$. During inference, the XObjAlign module and also exo mask prompt $P_{vis}$ are removed, and the remaining model is used for inference.
3.2 Multimodal Condition Fusion